If Pd = 100 - Qd and Ps = 0.25Qs
* determine the equilibrium quantity and price?
* Suppose the ad valorem tax of 10% is imposed on buyers of the wine,
graphically sketch the behavior of the demand and supply curves.
* what is the new equilibrium price and quantity?
* what is the price that buyers pay and supplies get ?
* Also determine the tax revenue and tax incidence on the buyers and sellers
1
Expert's answer
2021-02-09T07:04:12-0500
"Q_D=100-p"

"Q_S=4p"


"100-p=4p"

"p=20"

"Q=80"

After tax:


"100-p=3.6p"

"p=21.7"


"Q=78.3"